26. Exhibit 22-14 Quantity of Labor Output Marginal Physical Product (MPP) Wage Marginal Cost    0    ...

26.

Exhibit 22-14



Quantity of Labor



Output

Marginal Physical Product (MPP)



Wage



Marginal Cost

   0    

0

1

100

(C)

$600

(E)

        2        

(A)

75

600

(F)

      3      

(B)

50

600

(G)

          4          

240

(D)

600

(H)



Assume that labor is the only variable input and that each additional laborer is paid $600.

Refer to Exhibit 22-14. What is the MPP of the first unit of labor [blank (C)]?

a.

50 units

b.

100 units

c.

600 units

d.

700 units
Answer #1

ANswer
Option b
100 units

Marginal product of n th unit of labor =total product of n units of labor -total product of n-1 units of labor
MP(n)=TP(n)-TP(n-1)
MP(1)=100-0=100 units